OpenStudy (anonymous):

What Improper fraction falls in between 8/3 and 3

sammixboo (sammixboo):

Okay we need to change \(\sf \large \color{magenta}{\frac{8}{3}}\) into a mixed fraction (when we are done we will change the answer to improper) \(\sf \large \color{magenta}{\frac{8}{3}}\) --- \(\sf \large \color{magenta}{2\frac{2}{3}}\) Now we need to multiply \(\sf \large \color{magenta}{2\frac{2}{3}}\) by two, because there is no fraction between \(\sf \large \color{magenta}{2\frac{2}{3}}\) and \(\sf \large \color{magenta}{3}\)... \(\sf \large \color{magenta}{2\frac{2 x 2}{3 x 2}}\) --- \(\sf \large \color{magenta}{2\frac{4}{6}}\) So now we find the fraction between \(\sf \large \color{magenta}{2\frac{4}{6}}\) and \(\sf \large \color{magenta}{3}\)

sammixboo (sammixboo):

The fractin between \(\sf \large \color{magenta}{2\frac{4}{6}}\) and \(\sf \large \color{magenta}{3}\) is \(\sf \large \color{magenta}{2\frac{5}{6}}\)... Do you know what \(\sf \large \color{magenta}{2\frac{5}{6}}\) is in improper form?

sammixboo (sammixboo):